A 450 kg compressor slides down a loading ramp that is 7.0 m long. Initially at rest, the compressor’s speed at the bottom of the ramp is 1.1 m/s. What is the magnitude of the net force acting on the compressor?
1
Expert's answer
2013-01-14T09:02:33-0500
From energy conservation law: all potential energy goes tokinetic, except work of external net force.

We can formulate it next way: T=W-A, t- kinetic, W-potential, A - work.
A=F*S
W=mgh
T=mv^2/2
So,
mv^2/2=mgh-f*s
mgh-mv^2/2=f*s
(m/s)(gh-v^2/2)=F
(450/7)(10*7-1.1^2/2)=F=4.371kN